# ConceptCaps - a Distilled Concept Dataset for Interpretability in Music Models

### v1

@@ Abstract
Concept-based interpretability methods like TCAV require clean, well-separated positive and negative examples for each concept. Existing music datasets lack this structure: tags are sparse, noisy, or ill-defined. We introduce ConceptCaps, a dataset of 23k music-caption-audio triplets with explicit labels from a 200-attribute taxonomy. Our pipeline separates semantic modeling from text generation: a VAE learns plausible attribute co-occurrence patterns, a fine-tuned LLM converts attribute lists into professional descriptions, and MusicGen synthesizes corresponding audio. This separation improves coherence and controllability over end-to-end approaches. We validate the dataset through audio-text alignment (CLAP), linguistic quality metrics (BERTScore, MAUVE), and TCAV analysis confirming that concept probes recover musically meaningful patterns. Dataset and code are available online.
@@ 1.1 Main Contributions
    We introduce ConceptCaps, a large-scale, copyright-free dataset of 23k music-description pairs with rich, validated concept labels and matched counterexamples, specifically designed for concept-based explainability research.
@@ 3.6 Implementation
Source code, datasets and pre-trained models are publicly available online for reproducibility purposes [^1] https://anonymized-for-blind-review.com. We are using $\beta$-VAE Higgins et al. (2017) variant of variational auto-encoder with additional $\beta$ parameter, which controls KL divergence loss. From our experiments we concluded that $\beta=0.25$ yielded the best overall results. It suggests that minimizing reconstruction loss is more important in our case than mapping latent space to normal distribution. Higher $\beta$ values resulted in model collapse suggested by VAE predicting only the few most popular attribute combinations. VAE is implemented with a standard fully-connected architecture: encoder with DD-1024 hidden units, where $D$ is denotes our aspect taxonomy dimension, mapping to a 256-dimensional latent space, and a symmetric decoder. We use Adam optimizer Kingma (2014) with learning rate $6\times 10^{-5}$ and train for 150 epochs on the 1,890 distilled samples.
@@
Stage 2: LLM BERTScore-F1 Zhang et al. (2020) $\uparrow$ 0.9132 Contextual semantic similarity using BERT embeddings between generated and ground-truth captions.
BLEU Papineni et al. (2002) $\uparrow$ 0.2465 Unigram, bigram, trigram, and 4-gram overlap with reference captions.
ROUGE-L Lin (2004) $\uparrow$ 0.4263 Longest common subsequence F-score between generated and reference captions.
MAUVE Pillutla et al. (2021) $\uparrow$ 0.9236 Distribution divergence metric between generated and human-written caption distributions.
Stage 3: Audio CLAP Elizalde et al. (2023) $\uparrow$ 0.4745 Similarity between textual description and audio representation calculated using pre-trained CLAP model.
FAD Kilgour et al. (2019) $\downarrow$ 0.4603 Fréchet Audio Distance, quantifying divergence between real and generated audio distributions using CLAP embeddings.
@@ 5.4.2 Results
The results presented in Figure 6 confirm that the classifier has learned domain-appropriate musical concepts. The TCAV scores reveal clear variations in concept importance that align with musicological expectations. Notably, the importance of ”fast tempo” and ”slow tempo” contradict each other, across genres: while Jazz shows a high sensitivity to ”slow tempo” ($\sim$0.8) and low sensitivity to ”fast tempo” ($\sim$0.2), this relationship is inverted for Metal and Rock, where ”slow tempo” provides zero discriminative utility.
Furthermore, instrumentation concepts show high class-conditional relevance; for example, ”piano” and ”acoustic guitar” reach their peak importance within the Jazz class ($1.0$ and $\sim$0.9, respectively), while remaining negligible for high-energy genres like Metal. These findings confirm that our distilled dataset provides the precise, semantically coherent examples required to extract and audit specific musical concepts in downstream models.
@@ 6 Conclusions
We present ConceptCaps, a dataset of 23k music-description pairs with validated concept labels, built through a two-stage pipeline that separates semantic modeling (VAE) from linguistic generation (fine-tuned LLM). This separation yields measurable gains in both caption quality and audio-text alignment over monolithic approaches like LP-MusicCaps or zero-shot generation. Our TCAV experiments confirm that the dataset enables meaningful concept-based probing of music classifiers. While extending the taxonomy to non-Western musical traditions remains a next step, we hope ConceptCaps will facilitate further research in interpretability of music models.

### v2

# ConceptCaps: a Distilled Concept Dataset for Interpretability in Music Models

@@ Abstract
Concept-based interpretability methods like TCAV require clean, well-separated positive and negative examples for each concept. Existing music datasets lack this structure: tags are sparse, noisy, or ill-defined. We introduce ConceptCaps, a dataset of 21k music-caption-audio triplets with explicit labels from a 200-attribute taxonomy. Our pipeline separates semantic modeling from text generation: a VAE learns plausible attribute co-occurrence patterns, a fine-tuned LLM converts attribute lists into professional descriptions, and MusicGen synthesizes corresponding audio. This separation improves coherence and controllability over end-to-end approaches. We validate the dataset through audio-text alignment (CLAP), linguistic quality metrics (BERTScore, MAUVE), and TCAV analysis confirming that concept probes recover musically meaningful patterns. Dataset and code are available online.
@@ 1.1 Main Contributions
    We introduce ConceptCaps, a large-scale, copyright-free dataset of 21k music-description pairs with rich, validated concept labels and matched counterexamples, specifically designed for concept-based explainability research.
@@ 3.6 Implementation
Source code, datasets and pre-trained models are publicly available online for reproducibility purposes [^1] https://github.com/BrunoSienkiewicz/ConceptCaps. We are using $\beta$-VAE Higgins et al. (2017) variant of variational auto-encoder with additional $\beta$ parameter, which controls KL divergence loss. From our experiments we concluded that $\beta=0.25$ yielded the best overall results. It suggests that minimizing reconstruction loss is more important in our case than mapping latent space to normal distribution. Higher $\beta$ values resulted in model collapse suggested by VAE predicting only the few most popular attribute combinations. VAE is implemented with a standard fully-connected architecture: encoder with DD-1024 hidden units, where $D$ is denotes our aspect taxonomy dimension, mapping to a 256-dimensional latent space, and a symmetric decoder. We use Adam optimizer Kingma (2014) with learning rate $6\times 10^{-5}$ and train for 150 epochs on the 1,890 distilled samples.
@@
Stage 2: LLM BERTScore-F1 Zhang et al. (2020) $\uparrow$ 0.8988 Contextual semantic similarity using BERT embeddings between generated and ground-truth captions.
BLEU Papineni et al. (2002) $\uparrow$ 0.2119 Unigram, bigram, trigram, and 4-gram overlap with reference captions.
ROUGE-L Lin (2004) $\uparrow$ 0.3906 Longest common subsequence F-score between generated and reference captions.
MAUVE Pillutla et al. (2021) $\uparrow$ 0.9503 Distribution divergence metric between generated and human-written caption distributions.
Stage 3: Audio CLAP Elizalde et al. (2023) $\uparrow$ 0.5563 Similarity between textual description and audio representation calculated using pre-trained CLAP model.
FAD Kilgour et al. (2019) $\downarrow$ 0.5135 Fréchet Audio Distance, quantifying divergence between real and generated audio distributions using CLAP embeddings.
@@ 5.4.2 Results
The results presented in Figure 6 confirm that the classifier has learned domain-appropriate musical concepts. The TCAV scores reveal clear variations in concept importance that align with musicological expectations. Notably, the importance of ”slow tempo” varies significantly across genres, being highly relevant for Classical music ($\sim$0.8) while providing low relevance for genres like Metal.
Furthermore, instrumentation concepts show high class-conditional relevance; for example, ”acoustic guitar” reaches high importance within the Country class ($\sim$0.9), which aligns with its characteristic sound. Conversely, ”electric guitar” exhibit inverse relevance patterns, being highly important for Rock and Metal while irrelevant for Classical music.
With this analysis, we demonstrate that the classifier’s internal representations are meaningfully structured around musically relevant concepts, validating the effectiveness of our distilled dataset for concept-based interpretability.
@@ 6 Conclusions
We present ConceptCaps, a dataset of 21k music-description pairs with validated concept labels, built through a two-stage pipeline that separates semantic modeling (VAE) from linguistic generation (fine-tuned LLM). This separation yields measurable gains in both caption quality and audio-text alignment over monolithic approaches like LP-MusicCaps or zero-shot generation. Our TCAV experiments confirm that the dataset enables meaningful concept-based probing of music classifiers. While extending the taxonomy to non-Western musical traditions remains a next step, we hope ConceptCaps will facilitate further research in interpretability of music models.
@@ 6.1 Limitations
6.2 Future Work
We belive that this work introduces a strong foundation for future research in music dataset generation and interpretability. Having direct access to quality concept dataset will facilitate development of new XAI methods specifically tailored for music models, making them more transparent and trustworthy. This is especially important as music generation models become more prevalent in creative industries, raising concerns about authorship, originality, and ethical use.
While present in our dataset, aspects related to human voice are not represented in resulting audio due to additional requirement of providing lyrics and voice synthesis for each sample, which is out of scope for this thesis. We decided to still include such samples as they represent important musical concepts related to vocal music, singing styles, and lyrical themes. Future work could explore integrating lyrics generation to create vocal music samples, further enriching the dataset’s diversity.
Introducing human in the loop during the generation process could further enhance the quality of both captions and audio. For example, human reviewers could validate or refine generated captions before audio synthesis, ensuring higher semantic fidelity. Similarly, human evaluation of audio samples could help identify and filter out low-quality or misaligned generations. Incorporating human feedback loops would increase the dataset’s reliability for interpretability research.
Future work could also explore expanding the taxonomy of musical concepts used in the VAE. By providing source code along with models used in the generation pipeline, researchers can adapt the model to include more diverse or specialized musical attributes, such as cultural genres, production techniques, or emotional nuances. This would allow the creation of tailored datasets for specific research questions or applications, further enhancing the utility of the proposed approach.